# Multimodal Conversation Structure Understanding

### v2

@@ 2 Related work
Multi-party conversation understanding (i.e., dialogues with more than two participants) can be summarized as solving “who says what to whom” [6]. This work tends to consider a limited set of modalities, including text- [7, 8, 9] or audio-only [10, 11] approaches. Representative tasks include addressee recognition and conversation disentanglement; both require understanding the relationship between speakers—or more generally, conversation participants—and their utterances. They are designed and built upon text-based data, such as chatroom logs [12, 13]. Works involving audio-visual data focus on speaker activity, identity, and visual co-occurrence [14, 15, 16] and their long-range dependencies [17, 18], mostly without resolving the semantics of conversation structure, until more recently [19]. In this work, we believe that multiple modalities are useful in capturing different aspects of the conversation structure, and a robust understanding of conversational dynamics should mirror human understanding of social interactions.
@@
Video understanding involves reasoning over an extended temporal context that includes a sequence of images and/or audio signals/ It is fundamental to tasks ranging from movie genre classification to scene graph creation [20, 21, 22, 23, 24, 25, 26]. Such work focuses on multimodal representations of video amd parsing interactions between visible objects, exemplified by visual commonsense reasoning (VCR, 27) and TVQA [28, 29], require models to infer plausible interactions between characters in an image (or a sequence of frames sampled from videos) and reason over their actions. However, they typically do not explicitly resolve the underlying conversation structure and often discards speech data (audio or transcription). Those works converge in recent advances of foundational models: they are natively multimodal [30, 31, 32, 33], general-purpose models that can also reason over videos. In this work, we assess the capabilities of such models to resolve the conversation structure of such long-form videos. Related work that bears most resemblance to this study does consider multiple modalities for solving conversation structure (addressee recognition) but involves a controlled setting [34, 35]. This work represents interactions and settings represented on screen, and we discuss the impact of this in §6.
3 The TV-MMPC dataset
In this section, we describe how we construct our TV-MMPC (Multi-modal Multi-Party Conversation) dataset. We start from TVQA [28], which includes sampled video frames, audio of the original clips (60–90 seconds), and the associated subtitles. For annotation, we randomly sample 50 clips from the following shows in TVQA: The Big Bang Theory, Friends, House M. D., and How I Met Your Mother. To adapt TVQA for our purpose, we take the following steps:
Re-transcribe the audio on the sentence-level.
The original subtitles lack clear sentence boundaries, which makes them unsuitable for annotation aimed at understanding conversational structure. To address this, we transcribe the raw audio using Whisper [36] and align the transcription with the original subtitle data from TVQA.
Infer and standardize speaker labels.
We are interested in conversational participants, which, in the context of TV series, are all characters that appear in the clip. For this reason, we focus on post-processing Some of the subtitles from TVQA contain speaker labels; those labels are usually the first names of the characters in the clip. To facilitate downstream face recognition and annotation, we standardize those labels and map them to actor names by querying TMDb [^1] https://www.themoviedb.org/ for the cast list of each episode. This gives us the mapping between the canonical names of the characters and the actors who play them. With this information, we project TVQA speaker labels where possible (i.e., when the label is present and not unknown for the given sentence) and the algorithm described in [20, p. 24] to align TVQA subtitles and Whisper transcriptions. For a second pass, we infer speaker labels from the sampled frames and audio: we use the most frequently occurring face during the speech in question,  and we recognize the face following the pipeline described in [37], which maps the faces to actor names on IMDb, [^2] https://www.imdb.com/ and the character who dominates the duration of the sentence is treated as the speaker of the sentence. Those speaker labels are verified during the annotation stage; if no face or actor is detected from the automatic pipeline described above, we manually annotate them too.
Annotation guidelines and procedure.
For ease of annotation, we stitch the sampled frames with face captions and use the audio to create a rough approximation of the original clip. Annotators are carried out by co-authors of the paper; we watch the clips on a dedicated annotation interface and annotate conversational roles and reply-to for each utterance. For more details, see Supplementary Information.
Dataset statistics.
See Table 1.
Per clip Overall
# Unique # Avg. Total # Unique # Total
Speakers 4.18 21.89 156 4,378
Addressees 4.48 28.00 168 5,599
Side Participants 2.90 17.06 121 3,412
4 Conversation structure understanding
On a conceptual level, our goal is to answer the following questions for any video clip: a.) Who gets to start a conversation and hold the floor? b.) Who is being directly addressed? c.) Who listens in on the side? In articulating our proposed framework of multimodal conversation structure understanding, we build on two existing lines of work:
Conversation disentanglement.
Here, we want to resolve the conversation structure and detect the start of individual sub-conversations or conversational threads [38, 39]. We follow the formulation in [40]: Given an utterance of interest (UOI) $u_{t}$, the task is to identify its parent utterance $u_{p}$, which is the utterance that $u_{t}$ directly replies to ($u_{t}\rightarrow u_{p}$). In other words, we find the most relevant preceding utterance in the conversational thread; if the utterance marks the start of a new thread, we say it replies to itself. The thread is the transitive closure of such pairwise links; the main task for conversation disentanglement is, then, the reply-to task.
Conversational role attribution.
We draw on the standard tasks of speaker and addressee recognition [41, 42] and describe a more general task of conversational role attribution. According to [43], conversational roles are observably cast or projected by the speaker $s_{t}$ at time $t$, regardless of what happens at $t+1$ (e.g. addressee might not hear or understand the speaker, and that does not impact our annotation). We revise the taxonomy proposed in [44, 45] (see Table 2) by organizing participant roles along three binary dimensions: whether a participant is addressed by the speaker, ratified as a member of the conversational group, and known to be attending or perceiving the utterance. This reformulation clarifies distinctions among roles and highlights their dependency: addressees are explicitly addressed, ratified as participants, and known to be listening; side-participants are not addressed but still ratified and perceptually engaged. Bystanders, by contrast, are neither addressed nor ratified, and may not be known to be present, which, unlike in [45], includes non-active participants who are overhearers.
Table 2: Participant role matrix.
Role addressed ratified known
Addressee $+$ $+$ $+$
Side-participants -- $+$ $+$
Bystanders -- -- $\pm$
Formally, given an input video clip $V$, we first extract the sequence of utterances $\mathcal{U}=\{u_{1},u_{2},\dots,u_{|\mathcal{U}|}\}$, where each utterance $u_{i}\in\mathcal{U}$ corresponds to a segment of the video defined by a start time $t_{s,i}$ and an end time $t_{e,i}$, $t_{e,i}\leq t_{s,i+1}$. We also identify the set of participants $\mathcal{P}=\{p_{1},p_{2},\dots,p_{|\mathcal{P}|}\}$ appearing in the video, which are derived from face recognition or other identity tracking methods applied to $V$. For each utterance of interest (UOI) $u_{i}\in\mathcal{U}$, the goal is to predict its conversational context by identifying its speaker, addressees, side-participants, and the utterance it directly replies-to. A multi-party conversation structure solver $\mathcal{G}$ maps the current utterance and relevant context to these structural elements:
$\mathcal{G}:(u_{i},\mathcal{U},\mathcal{P})\mapsto\{u_{p},S_{i},\mathcal{A}_{i},\mathcal{E}_{i}\}$ (1)
where reply-to utterance $u_{p}\in\{u_{1},\dots,u_{i-1}\}$ is the parent utterance that $u_{i}$ directly replies to, identified by a reply-to function $R:u_{i}\mapsto u_{p}$ such that $p<i$. If $u_{i}$ initiates a new thread, $u_{p}=u_{i}$; speaker $S_{i}\in\mathcal{P}$ is the participant who produced utterance $u_{t}$; addressees $\mathcal{A}_{i}\subseteq\mathcal{P}$ is the set of participants primarily addressed by $S_{i}$ during $u_{i}$; side-participants $\mathcal{E}_{i}\subseteq\mathcal{P}$ is the set of participants actively listening or engaged in the conversation involving $u_{i}$, but not directly addressed by $S_{i}$. Note that $\mathcal{A}_{t}\cap\mathcal{E}_{t}=\varnothing$. Given this formulation, the bystanders are participants who are neither addressees nor side-participants but belonging to the same thread and can therefore be determined heuristically.
5 Experiments
In this section, we evaluate six popular audio-visual models and vision–language models on TV-MMPC (§5.3) and explore resource-efficient approaches for task adaptation (§5.4).
5.1 Evaluation metrics
@@ 5.1.1 Conversational roles
[…]
@@
5.1.2 Conversational thread
Reply-to involves first identifying the pairwise links between an utterance of interest and its parent utterance (child utterance replies to the parent), and secondly turning those links into conversational threads. We can see the former as equivalent to binary classification (i.e. whether two utterances exhibit the reply-to relationship) and the second as a clustering problem. The former is evaluated using the F1 score; the latter on the cluster-based metrics below:
@@ 5.1.2 Conversational thread
This metric measures the dissimilarity between clustering assignments, based on entropy and mutual information:
$\text{VI}(C,C^{\prime})=H(C)+H(C^{\prime})-2I(C,C^{\prime}),\quad\text{NVI}(C,C^{\prime})=1-\frac{\text{VI}(C,C^{\prime})}{\log_{2}N}$ (2)
@@
One-to-One Overlap (1–1, [47]).
This metric finds a one-to-one mapping between gold and predicted clusters that maximizes their overlap. It treats the contingency matrix $M$, where $M_{i,j}$ is the number of shared utterances between gold cluster $i$ and predicted cluster $j$, as a weighted bipartite graph. The optimal assignment is solved using a min-cost flow algorithm, which identifies the pairing of clusters that yields the greatest total intersection:
[…]
@@
This metric computes precision, recall, and F1-score based on exact identity between individual gold-standard clusters $c_{i}\in C_{\text{gold}}$ and predicted clusters $c^{\prime}_{j}\in C_{\text{pred}}$. A match is identified for a gold cluster $c_{i}$ if there exists any predicted cluster $c^{\prime}_{j}$ such that $c_{i}=c^{\prime}_{j}$ (i.e., they contain the identical set of elements). Letting $N_{\text{match}}$ be the total number of unique gold clusters $c_{i}$ for which such a match exists, Precision is defined as $P=N_{\text{match}}/|C_{\text{pred}}|$, Recall as $R=N_{\text{match}}/|C_{\text{gold}}|$, and F1 is their harmonic mean, measuring the perfect recovery of clusters.
5.2 Heuristic baseline
We first establish a heuristic baseline. We use Whisper [36] and pyannote [49] to obtain word-level timestamps and speaker labels. Since pyannote only provides generic speaker labels, we determine the speaker by aggregating the frequency of each face appearing at the word level, then assigning the speaker role to the face that appears most frequently within each sentence. For each utterance $u_{i}$, we consider faces present in the context window $[i{-}1,i]$. Among these, the most frequently occurring face excluding the speaker is designated as the addressee, while all remaining faces are labeled as side-participants. For the linking task, the previous utterance is treated as the parent utterance (i.e., $u_{i}$ always reples to $u_{i-1}$).
5.3 LLM zero-shot performance
We evaluate the zero-shot performance of two types of models on our tasks: vision–language models (LLaMA 4 Scout, [^3] https://huggingface.co/meta-llama/Llama-4-Scout-17B-16E GPT-4.1, [^4] https://openai.com/index/gpt-4-1/ o4-mini, [^5] https://openai.com/index/introducing-o3-and-o4-mini/ and Gemini 2.0 Flash [^6] https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash) and audio-visual LLMs (Reka-Flash [32], Qwen 2.5 Omini [50], and Gemini 2.0 Flash with reconstructed clips). The multimodal LLM $\mathcal{F}$ performs a structured prediction $o_{i}$ for each $u_{i}\in\mathcal{U}$:
where $C_{i}$ is the context of $u_{i}$, $\mathcal{P}_{V_{i}}=\{p_{1},p_{2},\dots,p_{|\mathcal{P}|}\}$ is the set of identified participants associated with $C_{i}$, which corresponds to the cast list of the episode from which the clip is taken. Participants include the text of character names and, where available, the best face crop between $t_{s,i}$ and $t_{e,i}$. $Q_{u_{i}}$ is the question associated with the $u_{i}$ in: it includes the line index $i$, the start and end timestamps $t_{s,i}$ and $t_{e,i}$, and the question text itself, which prompts the model for conversation structure analysis. The output ${o}_{i}$ contains the index of the reply-to utterance $o_{r_{i}}\leq i$, identities of the predicted speaker $s_{i}$, addressees $a_{i}$, side-participants $e_{i}$; note $o_{a_{i}},o_{e_{i}}\subseteq\mathcal{P}_{V_{i}}$. Given the design of models, we need to structure the overall prompt differently for each class of models, particularly what context information $C_{i}$ to put in the model:
Vision–language models.
The standard setup in video understanding often involves sampling visual frames from a video while discarding its audio. Here, context $C_{i}=\mathcal{U}^{\star}_{i}$, and $\mathcal{U}_{i}=\{(v^{\star}_{j},u_{j},\;f_{j})\}_{j=1}^{|\mathcal{U}|}$: $v^{\star}_{i}$ is the final frame sampled from $V_{i}$ in the interval $[t_{s,j}\,,t_{e,j}]$, interleaving with the words of $u_{j}$; this is further augmented with audio-derived speaker features $f$ extracted using Librosa [51]: speech rate (slower or faster), pitch range (wide or narrow), pitch mean (lower or higher), and spectral centroid (darker or brighter), which collectively aim to capture the acoustic profile of the speaker.
Audio-visual LLMs.
The results are reported in Table 3. All vision–language models we consider perform significantly better than the heuristic baseline across all metrics, other than on the reply-to task. Another exception is LLaMA 4 Scout, which falls short on side-participant attribution, although it achieves better results on utterance linking than other models, with a statistically significant margin over GPT-4.1 mini. Among the audio-visual language models we consider, Gemini 2.0 Flash outperforms others by a considerable margin; Qwen 2.5-Omni 7B and Reka-Flash do not show statistically significant improvements over the heuristic baseline on most metrics. These results suggest the need for further research into multimodal models, particularly in developing methods that more effectively leverage and integrate different input modalities.
5.4 Resource-efficient supervised fine-tuning
Table 4: Performance of LoRA SFT of Qwen 2.5-Omni 7B compared to zero-shot. Boldface indicates statistically significant differences, determined by non-overlapping 95% CIs.
[…]
Given the increasing popularity of audio-visual LLMs and the growing number of open-source implementations in this space, we turn to explore task-specific adaptation through supervised fine-tuning (SFT). In this subsection, we describe our approach using Low-Rank Adaptation (LoRA) to fine-tune Qwen 2.5-Omni 7B within the LLaMA Factory framework [52]. To minimize computational overhead, LoRA is applied across all layers of the model with a lightweight rank of 2, and a total training steps of 3,125 on one L40S GPU. Each training sample is limited to 4,096 tokens to manage memory and processing time, and the total number of random samples is capped at 1,000. We employ a leave-one-show-out evaluation strategy: training the model on all but one show, and testing it on the held-out show. This process is repeated for each show in the dataset. The results are reported in Table 3, which we compare the performance before and after LoRA SFT on the same TV series. After fine-tuning, despite the constraint in resources, the model has a statistically significant boost on both addressee and side-participant recognition.
6 Discussion
In this section, we examine key factors that influence model performance of the most capable model from the previous section, Gemini 2.0 Flash. We first analyze the potential impact of participant anonymization and the presence of TV-specific dialogue patterns (§6.1), and then turn to how different modalities of input contribute to or limit the model’s effectiveness across tasks (§6.2).
6.1 Effects of anonymization and non-dialogic lines
Table 5: Impact of anonymization and non-dialogic lines on the performance Gemini 2.0 Flash on conversation structure understanding. Boldface indicates significance by non-overlapping 95% CIs.
[…]
Prior work has shown that anonymizing speakers can influence model learning and downstream performance in text-based dialogue understanding tasks [53, 54]. Motivated by this, we investigate whether anonymizing the identities of conversational participants has a similar impact in our multimodal setting. To do so, for each clip, we create an anonymization mapping based on its individual episode cast list: we assign a random letter to each character within a clip (e.g., “Sheldon Cooper” becomes Character C). We then apply the same frame captioning procedure described in §3, but we replace those actual character names with their anonymized labels. In prompting Gemini, our setup is the same as described in §5.3: the model sees the reconstructed video clips, along with the face crop of the highest confidence score. The difference in performance is presented in Table 5. We observe significant drops in the precision and F1 scores for speaker and addressee recognition, by 64.92 and 52.37 points, respectively.
This points to future work in two directions: first, improving identity-awareness in video understanding [55];  second, assessing the degree to which the performance drop can be attributed to the phenomenon of memorization in the audio-visual LLMs, which is relevant in the context of vision–language models [56] and ASR models [57]. This is more than a straightforward case of test set contamination, where test data overlaps with training data, but a subtler one, where models would leverage their parametric knowledge, among other things, acquired from exposure to cultural artifacts and information about them, potentially across different modalities, when solving a task. Such memorization can impact downstream performance and raises further questions about cultural representation, which has been explored in the context of literary texts [58, 59] and can be relevant to audio-visual LLMs.
Since our dataset is derived from TVQA, during our annotation process, we note that there are edge cases in conversational role attribution that are specific to the medium of television. For example, there are narrators that fill in the background of the story, episode recap segments, or expository monologues for characterization: none of them are part of the actual conversation that takes place in the clip. We annotate those utterances as extra-diegetic (spoken by narrators or characters outside the immediate scene) and monologue (extended speech directed at no one in particular), collectively referred to as non-dialogic lines. While we expect those lines present edge cases and can be challenging to the model, Table 5 (second row) suggests that removing those utterances in evaluation does not lead to significant improvement on those metrics.
6.2 Effects of different modalities of input
Figure 2: Signed explained rank variance from Spearman’s $\rho$ between clip-level features and F1 scores for individual conversational roles. Bars indicate direction and magnitude of correlation: solid ones are significant ($p<0.05$); hatched ones are not. Four categories of features are extracted and presented: those related to length, textual diversity, acoustic clarity, and visual face presence.
To better understand the factors driving performance variation in conversational role attribution, we examine the correlations between interpretable, clip-level features (across textual, acoustic, and visual modalities) and F1 scores (across three roles). We employ Spearman’s $\rho$, a non-parametric measure without linear or normal assumptions, to assess monotonic relationships between ranked variables: a positive $\rho$ indicates that higher values of a feature consistently co-occur with higher ranks of F1 scores, while a negative one suggests higher feature values accompany lower F1 performance.
We present those relationships in Figure 2: for ease of presentation, the impact of each feature is depicted in the signed explained rank-variance: $R^{2}=\text{sign}(\rho)\cdot\rho^{2}\times 100$. Statistically significant correlations ($p<0.05$) are distinguished by bars with solid color fill, while non-significant relationships appear in diagonally hatched bars. It is not surprising that of the features we consider, the number of conversational participants has the strongest negative influence, substantially reducing F1 scores across all roles, which accounts for approximately 3–8% of variance. The influence of clip length is negligible, primarily because TVQA has clips of mostly 60–90 seconds in length.
For textual features, we consider mean sentence length (measured by tokens separated by whitespace) and type–token ratio. We see no significant correlation for this category of features, which suggests that lexical diversity alone does not meaningfully impact model performance. Acoustic features (mean pitch and spectral centroid), indicative of speech clarity and brightness, exert small yet consistent positive correlations with performance, with significance achieved primarily for side-participant attribution (pitch: $\rho=+0.17$, $p=0.018$; spectral centroid: $\rho=+0.15$, $p=0.033$), which suggests that clearer, acoustically brighter speech helps in identifying those peripheral conversational participants. For visual features, we consider the ratio of detected faces relative to the number of unique conversational participants present in annotated data (max face coverage) and raw detected face counts, both capture representational fidelity in the clip. Here we see role-specific dynamics: max face coverage positively correlates with speaker role attribution ( $\rho=+0.15$, $p=0.044$), which sugests accurate face detection can benefit this task. Raw face counts, on the other hand, negatively correlate with side-participant detection ($\rho=-0.23$, $p=0.0012$), which confirms that visual clutter degrades the model performance on this task.
Overall, we observe that features from different modalities contribute differently to role attribution performance: face crowding and acoustic clarity exhibit consistent effects on performance, while textual features offer weaker signals. This suggests that conversation structure understanding is not purely a textual problem: future work on this task might benefit from looking further into grounding in the perceptual context in which conversation unfolds. For future development of audio-visual LLMs, cross-modal cues (e.g. face–participant identity alignment, prosodic salience, etc.) have the potential to enable better parsing of conversation structure.
7 Conclusion
Conversations are usually structured by roles—who is speaking, who’s being addressed, and who’s listening—and unfold in threads that break with changes in speaker floor or topical focus. While large language models (LLMs) are increasingly proficient at dialogue and reasoning, evaluation of their performance on fine-grained conversation structure understanding in multimodal, multi-party settings remains relatively limited. This work is our attempt to bridge this gap and introduce tasks targeting conversational role attribution and conversation disentanglement. To support future work, we present TV-MMPC, a human-annotated dataset built on TVQA, with 4,378 speaker and reply-to annotations, 5,599 adrssee labels, and 3,412 side-participant tags.
We evaluate popular audio-visual LLMs and vision–language models on our dataset, and our experimental results suggest that multimodal conversational structure understanding remains challenging. The most performant audio-visual LLM, Gemini 2.0 Flash, outperforms all vision–language models across all metrics, especially in speaker and addressee recognition. However, its performance drops significantly when conversation participants are anonymized, which points to identity-awareness and memorization in audio-visual LLMs as interesting directions for further research, especially in relation to cultural representation. Additional analysis on Gemini suggests that the number of conversation participants in a clip is the strongest negative predictor of role-attribution performance, while acoustic clarity and detected face coverage yield positive associations. On the other hand, performance of the open-source Qwen 2.5-Omini 7B demonstrates improved performance even with low-rank fine-tuning in a resource-constrained setup, using only a subset of our annotated data. We hope this work lays the groundwork for future evaluation and development of multimodal LLMs that can reason more effectively about conversation structure.
8 Limitations
This work introduces a complex set of tasks that required developing detailed annotation guidelines and extensive annotators, and ensuring quality entails a time-consuming and labor-intensive process. We leave large-scale crowdsourcing, model distillation, or other approaches to scale up annotation, which this work enables along with the annotation guidelines, for future work. To maximize representativeness, we sample clips from TVQA entirely at random. This is not exhaustive of the original dataset and may overlook certain conversational patterns. Relatedly, the TV shows represented in this work is limited to the English language and American culture. We rely extensively on metadata from TMDb and IMDb for actor and character identification; while we believe they are generally reliable and accurate, these sources can include uncredited characters or omit minor roles. The effectiveness of our TVQA processing pipeline is limited by the capabilities of ASR and face recognition models, which is beyond the scope of this work.
Acknowledgments and Disclosure of Funding
Appendix A Supplementary information
A.1 Annotation guidelines

### v3

@@ Appendix B Additional related work
Multi-party conversation understanding (i.e., dialogues with more than two participants) can be summarized as solving “who says what to whom” Gu et al. (2022). This work tends to consider a limited set of modalities, including text- (Lu et al., 2022; Tan et al., 2023; Penzo et al., 2024) or audio-only (Akhtiamov et al., 2019; Lerner et al., 2022) approaches.
Representative tasks include addressee recognition and conversation disentanglement; both require understanding the relationship between speakers—or more generally, conversational participants—and their utterances. They are designed and built upon text-based data, such as chatroom logs Kummerfeld et al. (2019); Gu et al. (2021). Works involving audio-visual data focus on speaker activity, identity, and visual co-occurrence Bredin and Gelly (2016); Roth et al. (2020); Sharma and Narayanan (2023) and their long-range dependencies Tao et al. (2021); Kim et al. (2024), mostly without resolving the semantics of conversation structure, until more recently Korbar et al. (2024).
Text–vision models leverage image captions or scene descriptions to infer conversational roles in the broader video understanding framework, which suggests a focus on reasoning over the visual aspect of people and objects in the frame that ignores speech (Zellers et al., 2019; Lei et al., 2020). In this work, we believe that multiple modalities are useful in capturing different aspects of the conversation structure, and a robust understanding of conversational dynamics should mirror human understanding of social interactions.
@@
Video understanding involves reasoning over an extended temporal context that includes a sequence of images and/or audio signals. It is fundamental to tasks ranging from movie genre classification to scene graph creation Huang et al. (2020); Vicol et al. (2018); Wu and Krahenbuhl (2021); Islam and Bertasius (2022); Mangalam et al. (2023); Zhang et al. (2023); Sun et al. (2024). Such work focuses on multimodal representations of video and parsing interactions between visible objects, exemplified by visual commonsense reasoning (VCR, Zellers et al., 2019) and TVQA (Lei et al., 2018, 2020), require models to infer plausible interactions between characters in an image (or a sequence of frames sampled from videos) and reason over their actions. However, they typically do not explicitly resolve the underlying conversation structure and often discards speech data (audio or transcription).
Those works converge in recent advances of foundational models: they are natively multimodal Gemini Team et al. (2023); OpenAI et al. (2024); Reka Team et al. (2024); AI@Meta (2025), general-purpose models that can also reason over videos. In this work, we assess the capabilities of such models to resolve the conversation structure of such long-form videos. Related work that bears most resemblance to this study does consider multiple modalities for solving conversation structure (addressee recognition) but involves a controlled setting (Le Minh et al., 2018; Inoue et al., 2025). This work represents interactions and settings represented on screen, and we discuss the impact of this in §E.
Appendix C Evaluation metrics
@@ C.1 Conversational roles
Speaker identification is evaluated by comparing the predicted speaker label with the true label: $\nicefrac{{1}}{{N}}\sum_{i=1}^{N}\text{1}(y_{i}=\hat{y}_{i})$, where $y_{i}$ is the true label, $\hat{y}_{i}$ is the predicted label. This is equivalent to precision @1 used in previous work Gu et al. (2023).
@@
C.2 Conversational thread
Reply-to involves first identifying the pairwise links between an utterance of interest and its parent utterance (child utterance replies to the parent), and secondly turning those links into conversational threads. We can see the former as equivalent to binary classification (i.e., whether two utterances exhibit the reply-to relationship) and the second as a clustering problem. The former is evaluated using the F1 score; the latter on the cluster-based metrics below:
@@ C.2 Conversational thread
This metric measures the agreement between two sets of clusters. Based on information theory, it asks, “How much information does the predicted clustering give me about the actual (gold) clustering?” A higher score indicates that the predicted and gold threads are more structurally similar. Formally, this metric measures the dissimilarity between clustering assignments $C$ and $C^{\prime}$, based on entropy and mutual information:
$\text{VI}(C,C^{\prime})$ $=H(C)+H(C^{\prime})-2I(C,C^{\prime})$ (4)
$\text{NVI}(C,C^{\prime})$ $=1-\frac{\text{VI}(C,C^{\prime})}{\log_{2}N},$ (5)
@@
One-to-One Overlap (1–1, Elsner and Charniak, 2008).
This is a more forgiving metric that rewards identifying the essence of each conversational thread. It finds the best possible one-to-one pairing between the predicted and gold clusters to maximize their shared utterances. It doesn’t require a perfect match, so it gives credit even if a predicted thread is mostly correct but has a few extra or missing utterances. Formally, this metric finds a one-to-one mapping between gold and predicted clusters that maximizes their overlap. It treats the contingency matrix $M$, where $M_{i,j}$ is the number of shared utterances between gold cluster $i$ and predicted cluster $j$, as a weighted bipartite graph. The optimal assignment is solved using a min-cost flow algorithm, which identifies the pairing of clusters that yields the greatest total intersection:
$\text{One-to-One Overlap}=\frac{\max_{\pi}\sum_{i}M_{i,\pi(i)}}{\sum_{i,j}M_{i,j}},$ (6)
@@
This is our strictest metric; it measures the perfect recovery of threads. A predicted cluster only gets credit if it is an identical match to a gold-standard cluster (containing the exact same set of utterances). This score reflects the model’s ability to precisely identify the complete boundaries of a conversation. Formally, this metric computes precision, recall, and F1-score based on exact identity between individual gold-standard clusters $c_{i}\in C_{\text{gold}}$ and predicted clusters $c^{\prime}_{j}\in C_{\text{pred}}$. A match is identified for a gold cluster $c_{i}$ if there exists any predicted cluster $c^{\prime}_{j}$ such that $c_{i}=c^{\prime}_{j}$ (i.e., they contain the identical set of elements). Letting $N_{\text{match}}$ be the total number of unique gold clusters $c_{i}$ for which such a match exists, Precision is defined as $P=N_{\text{match}}/|C_{\text{pred}}|$, Recall as $R=N_{\text{match}}/|C_{\text{gold}}|$, and F1 is their harmonic mean, measuring the perfect recovery of clusters.
Appendix D Additional experimental details
We do not change the default temperature of the LLMs that we evaluate. We use the Pydantic schema for structured output (see Fig. 6). Our system instructions for Gemini 2.0 Flash, similar to those used for other models we consider, are in Fig. 9. This represents a succinct summary of our annotation guidelines. For our LLaMA 4 Scout experiments, we run the model in full precision on four H100 GPUs served by vLLM Kwon et al. (2023), which takes around 3 hours to complete. For supervised fine-tuning with Qwen 2.5-Omni 7B, to minimize computational overhead, LoRA is applied across all layers of the model with a lightweight rank of 2, and a total training steps of 3,125 on one L40S GPU. Each training sample is limited to 4,096 tokens to manage memory and processing time, and the total number of random samples is capped at 1,000. Qwen 2.5-Omni 7B inference is performed with full precision with Flash Attention 2 enabled Dao (2024) on four L40S GPUs, which takes around 7 hours. For Qwen 2.5-Omni 7B LoRA SFT, we follow the instructions and configurations provided on https://github.com/hiyouga/LLaMA-Factory/pull/7537.
class ConversationalRoles(BaseModel):
line_index: int
reply_to: int
speaker: str
addressees: list[str]
side_participants: list[str]
class ClipRoles(BaseModel):
clip_roles: list[ConversationalRoles]
Figure 6: Pydantic schema.
Appendix E Gemini error analysis
To further illustrate these quantitative findings, we discuss an edge-case clip that models perform badly on: House M.D., season 4, episode 2, segment 2, clip 11:
  • Visual clutter: The scene features a large number of participants, where a group of four or more doctors discusses a case in a busy hallway. This directly relates to our quantitative finding that participant count is the strongest negative predictor of performance across all conversational roles.
  • Visual-conversational misalignment: The editing in this scene breaks a simple alignment between sight and sound: for instance, the dialogue of the medical team is heard while the camera follows House, a different, lone character walking through another part of the hospital. Models that rely heavily on the visual presence of a speaker’s face to assign roles more frequently fail here, which supports our quantitative result that lower face coverage correlates with weaker speaker attribution performance.
  • Speech bleed-over: The audio from one conversation begins before the speakers are visually introduced, playing over shots of other scenes or characters. This editing technique, where audio precedes the corresponding video, creates significant ambiguity that makes it difficult for a model to correctly segment and attribute the initial utterances. This aligns with our finding that lower acoustic clarity can negatively impact role attribution.
Appendix F Annotation guidelines